# DrLLM: Prompt-Enhanced Distributed Denial-of-Service Resistance Method with Large Language Models

### v1

@@ Abstract
The increasing number of Distributed Denial of Service (DDoS) attacks poses a major threat to the Internet, highlighting the importance of DDoS mitigation. Most existing approaches require complex training methods to learn data features, which increases the complexity and generality of the application. In this paper, we propose DrLLM, which aims to mine anomalous traffic information in zero-shot scenarios through Large Language Models (LLMs). To bridge the gap between DrLLM and existing approaches, we embed the global and local information of the traffic data into the reasoning paradigm and design three modules, namely Knowledge Embedding, Token Embedding, and Progressive Role Reasoning, for data representation and reasoning. In addition we explore the generalization of prompt engineering in the cybersecurity domain to improve the classification capability of DrLLM. Our ablation experiments demonstrate the applicability of DrLLM in zero-shot scenarios and further demonstrate the potential of LLMs in the network domains. DrLLM implementation code has been open-sourced at https://github.com/liuup/DrLLM.
@@ II-1 Cyber IDS and Method
For DDoS traffic detection and classification tasks, existing IDSs are usually based on neural network architectures and the final results reached state-of-the-art (SOTA). For the problem of imbalanced data categories in the data set, most methods use upsampling or downsampling. Yisroel Mirsky et al.[12] presented Kitsune based on autoencoders to detect abnormal traffic data. Lucid[13] was a lightweight deep learning solution based on convolutional neural network (CNN) in recource-constrained environments. IDS-INT[14] using transformer-based transfer learning for imbalanced network traffic and the Synthetic Minority Oversampling Technique (SMOTE) is implemented to balance abnormal traffic and detect minority attacks. In practical scenarios, they are mostly deployed on the data plane or control plane of a programmable switch[15, 16] to handle large-scale hybrid and dynamic traffic attacks, and differentiate between normal and abnormal patterns.
@@ III Overview and Design
The Knowledge Embedding module is primarily tasked with the extraction of the global information of the dataset and their subsequent embedding into the Knowledge Prompt. The objective is to augment the LLMs’ expert knowledge, mitigate output hallucinations, and enhance classification accuracy. The Token Embedding module is primarily tasked with encoding network flow data, aligning tabular data to the text modality, and leveraging multiple technologies to enhance accuracy. The Role Reasoning module is responsible for integrating the aforementioned two modules, constructing the reasoning pipeline, and obtaining classification results through LLMs’ output. In general, our objective is to facilitate the analysis of network traffic data by LLMs through the construction of associations between global prior knowledge and local information, thereby enabling the accurate classification of data.
@@ IV-C Prompts and Evaluation Metrics
As mentioned in Section III-C, we use regular expressions to extract classification probabilities from the output and constrain them using CoD. However, during the evaluation process, we found that the LLMs occasionally outputs in an incorrect format. In order to verify the effectiveness of multiple modules of DrLLM, we carefully crafted four prompt templates $P_{0}=BP(BasicPrompt)$, $P_{1}=BP+CoD$, $P_{2}=BP+CoD+CoT$, $P_{3}=KP+BP+CoD+CoT$ for comparison. Due to space constraints, we will present the full prompt templates in the open source repository.
@@ IV-D Ablation Study
TABLE I: Accuracy of DrLLM with Different LLM Backbones
Backbones Metrics5 $P_{0}$ $P_{1}$ $P_{2}$ $P_{3}$
GPT1 F1 $0.6573$ $0.7024$ $0.7197$ $0.7732$
Recall $0.9033$ $0.6219$ $0.6620$ $0.7571$
AUC $0.7208$ $0.7797$ $0.7957$ $0.8345$
Llama32 F1 $0.5117$ $0.5275$ $0.2951$ $0.5904$
Recall $0.3567$ $0.3744$ $0.1737$ $0.4415$
AUC $0.6613$ $0.8543$ $0.7614$ $0.8124$
Deepseek3 F1 $0.7672$ $0.7904$ $0.8114$ $0.8499$
Recall $0.6507$ $0.6994$ $0.6882$ $0.8213$
AUC $0.8710$ $0.8853$ $0.8927$ $0.9212$
Qwen24 F1 $0.6525$ $0.6396$ $0.6968$ $0.7436$
Recall $0.8307$ $0.7448$ $0.7383$ $0.7555$
AUC $0.6984$ $0.7038$ $0.7584$ $0.8092$
1Denotes GPT-4o-mini. 2Denotes Llama3-70b.
3Denotes Deepseek-chat-v2. 4Denotes Qwen2-57b-a14b-instruct.
5The best value is in bold.
@@ IV-E2 Confidence Lost
Backbones Metrics1,2 $P_{0}$ $P_{1}$ $P_{2}$ $P_{3}$
GPT $L_{1}$ $0.8760$ $0.1807$ $0.0037$ $0.0014$
$L_{2}$ $0.0340$ $0.0302$ $0.0192$ $0.0153$
Llama3 $L_{1}$ $0.2829$ $0$ $0.0003$ $0.0010$
$L_{2}$ $0$ $0.0048$ $0$ $0.0010$
Deepseek $L_{1}$ $0.0309$ $0$ $0$ $0.0020$
$L_{2}$ $0.0707$ $0.0309$ $0.0146$ $0.0119$
Qwen2 $L_{1}$ $0.5346$ $0.0129$ $0.0065$ $0.0027$
$L_{2}$ $0.1403$ $0.0931$ $0.0245$ $0.0092$
1The unit of $L_{1}$ and $L_{2}$ is percentage. Lower is better.
2The best value is in bold.

### v3

@@ Abstract
The increasing number of Distributed Denial of Service (DDoS) attacks poses a major threat to the Internet, highlighting the importance of DDoS mitigation. Most existing approaches require complex training methods to learn data features, which increases the complexity and generality of the application. In this paper, we propose DrLLM, which aims to mine anomalous traffic information in zero-shot scenarios through Large Language Models (LLMs). To bridge the gap between DrLLM and existing approaches, we embed the global and local information of the traffic data into the reasoning paradigm and design three modules, namely Knowledge Embedding, Token Embedding, and Progressive Role Reasoning, for data representation and reasoning. In addition we explore the generalization of prompt engineering in the cybersecurity domain to improve the classification capability of DrLLM. Our ablation experiments demonstrate the applicability of DrLLM in zero-shot scenarios and further demonstrate the potential of LLMs in the network domains. The implementation of DrLLM is available at https://github.com/liuup/DrLLM.
@@ II-1 Cyber IDS and Method
For DDoS traffic detection and classification tasks, existing IDSs are usually based on neural network architectures and the final results reached state-of-the-art (SOTA). Yisroel Mirsky et al.[12] presented Kitsune based on autoencoders to detect abnormal traffic data. Lucid[13] was a lightweight deep learning solution in recource-constrained environments. IDS-INT[14] used transformer-based transfer learning for imbalanced network traffic and detect minority attacks. In practical scenarios, they are mostly deployed on the data plane or control plane of a programmable switch[15, 16] to handle large-scale hybrid and dynamic traffic attacks, and differentiate between normal and abnormal patterns.
@@ IV-A Datasets and Data Preprocessing
TABLE I: Accuracy of DrLLM with Different LLM Backbones
Backbones Metrics1 $P_{0}$ $P_{1}$ $P_{2}$ $P_{3}^{{}^{\prime}}$ $P_{3}$
GPT-4o-mini F1 $0.6573_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 14.99\%}}$ $0.7024_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 9.16\%}}$ $0.7197_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 6.92\%}}$ $0.7510_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 2.87\%}}$ $0.7732$
Recall $0.9033_{\color[rgb]{0,0.653,0}\mathbf{\blacktriangle 19.31\%}}$ $0.6219_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 17.86\%}}$ $0.6620_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 12.56\%}}$ $0.7308_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 3.47\%}}$ $0.7571$
AUC $0.7208_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 13.62\%}}$ $0.7797_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 6.57\%}}$ $0.7957_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 4.65\%}}$ $0.8254_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 1.09\%}}$ $0.8345$
Llama3-70b F1 $0.5117_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 13.33\%}}$ $0.5275_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 10.65\%}}$ $0.2951_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 50.02\%}}$ $0.5403_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 8.49\%}}$ $0.5904$
Recall $0.3567_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 19.21\%}}$ $0.3744_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 15.20\%}}$ $0.1737_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 60.66\%}}$ $0.5037_{\color[rgb]{0,0.653,0}\mathbf{\blacktriangle 14.09\%}}$ $0.4415$
AUC $0.6613_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 18.60\%}}$ $0.8543_{\color[rgb]{0,0.653,0}\mathbf{\blacktriangle 5.15\%}}$ $0.7614_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 6.28\%}}$ $0.6789_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 16.43\%}}$ $0.8124$
Qwen2-57b-a14b-instruct F1 $0.6525_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 12.25\%}}$ $0.6396_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 13.98\%}}$ $0.6968_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 6.29\%}}$ $0.7090_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 4.65\%}}$ $0.7436$
Recall $0.8307_{\color[rgb]{0,0.653,0}\mathbf{\blacktriangle 9.95\%}}$ $0.7448_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 1.41\%}}$ $0.7383_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 2.27\%}}$ $0.7554_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 0.01\%}}$ $0.7555$
AUC $0.6984_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 13.69\%}}$ $0.7038_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 13.02\%}}$ $0.7584_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 6.27\%}}$ $0.7310_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 9.66\%}}$ $0.8092$
Deepseek-chat-v2 F1 $0.7672_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 9.73\%}}$ $0.7904_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 7.00\%}}$ $0.8114_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 4.53\%}}$ $0.7759_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 8.71\%}}$ $0.8499$
Recall $0.6507_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 20.77\%}}$ $0.6994_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 14.84\%}}$ $0.6882_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 16.21\%}}$ $0.7256_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 11.65\%}}$ $0.8213$
AUC $0.8710_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 5.45\%}}$ $0.8853_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 3.90\%}}$ $0.8927_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 3.09\%}}$ $0.8787_{\color[rgb]{0.753,0,0}\mathbf{\blacktriangledown 4.61\%}}$ $0.9212$
1We highlight the $P_{3}$ by $\bullet$.
2If other prompts’ perform better than $P_{3}$ on a metric we denote it by $\bullet$, if not as well we denote it by $\bullet$.
@@ IV-C Prompts and Evaluation Metrics
As mentioned in Section III-C, we use regular expressions to extract classification probabilities from the output and constrain them using CoD. However, during the evaluation process, we found that the LLMs occasionally outputs in an incorrect format. In order to verify the effectiveness of multiple modules of DrLLM, we carefully crafted four prompt templates $P_{0}=BP(BasicPrompt)$, $P_{1}=BP+CoD$, $P_{2}=BP+CoD+CoT$, $P_{3}^{{}^{\prime}}=KP+BP+CoD$, $P_{3}=KP+BP+CoD+CoT$ for comparison. Due to space constraints, we will present the full prompt templates in the open source repository.
@@ IV-E2 Confidence Lost
Backbones Metrics5,6 $P_{0}$ $P_{1}$ $P_{2}$ $P_{3}^{{}^{\prime}}$ $P_{3}$
GPT1 $L_{1}$ $87.60$ $18.07$ $0.37$ $0.33$ $0.14$
$L_{2}$ $3.40$ $3.02$ $1.92$ $1.84$ $1.53$
Llama32 $L_{1}$ $28.29$ 00 $0.03$ $0.07$ $0.10$
$L_{2}$ 00 $0.48$ 00 $0.17$ $0.10$
Qwen23 $L_{1}$ $53.46$ $1.29$ $0.65$ $0.30$ $0.27$
$L_{2}$ $14.03$ $9.31$ $2.45$ $0.97$ $0.92$
Deepseek4 $L_{1}$ $3.09$ 00 00 $1.27$ $0.20$
$L_{2}$ $7.07$ $3.09$ $1.46$ $1.19$ $0.90$
1Denotes GPT-4o-mini. 2Denotes Llama3-70b.
3Qwen2-57b-a14b-instruct. 4Denotes Denotes Deepseek-chat-v2.
5We highlight the best value by $\bullet$.
6The unit of $L_{1}$ and $L_{2}$ is percentage [%]. Lower is better.